# Neural Network ANN Artificial Neural Network

### v1

@@ IV-B Online (real-time) experiment
In this experiment, we integrated the SNN based ball detection into our table tennis robot system, described in Section III-A. As explained in [2], to determine the 3D position of the table tennis ball, at least two cameras need to detect the ball in order to perform triangulation. Since we only have one device for each of the three SNN frameworks, covered in this work, we used one event-based camera with the introduced SNN ball detection together with a frame-based camera and the ball detection presented in [2]. We used the Akida PCIe card from BrainChip, since the system integration does not introduce latencies as high as the DynapCNN does, and we did not yet have access to a Loihi2 edge device. The experiment setup is depicted in Fig. 5 on the right, and the physical setup shown in Fig. 6.
Figure 6: The table tennis robot setup with the robot hitting back a table tennis ball in a rally.
In this experiment, we shot 15 table tennis balls with a Butterfly Amicus ball gun. Since the 3D triangulation and robot arm control are not part of this work, we used the existing ones from [2]. We report the ball return rate of the robot, since controlling the landing point of the ball is out of scope for this work. Given this setup, we achieved a ball return rate of $1.0$.
@@ V CONCLUSION
In this work, we used an event-based camera in combination with Spiking Neural Networks (SNNs) for ball detection in a table tennis robot system as a use case for a real-time perception task. Three different SNN frameworks, namely sinabs, MetaTF, and Lava, and a corresponding neuromorphic edge device (DynapCNN, Akida, and Loihi2) were used, and their errors, time per forward pass, and inference time were compared. Furthermore, we deployed one solution in an existing table tennis robot system to show the real-time capability of our proposed solution.
Despite the promise of asynchronous processing with SNNs, the current edge devices face limitations, primarily attributed to hardware integration. This could be solved by integrating a high-resolution event-based camera together with an SNN in one chip.

### v2

@@ IV-B Online (real-time) experiment
In this experiment, we integrated the SNN based ball detection into our table tennis robot system, described in Section III-A. As explained in [8], to determine the 3D position of the table tennis ball, at least two cameras need to detect the ball in order to perform triangulation. Since we only have one device for each of the three SNN frameworks, covered in this work, we used one event-based camera with the introduced SNN ball detection together with a frame-based camera and the ball detection presented in [8]. We used the Akida PCIe card from BrainChip, since the system integration does not introduce latencies as high as the DynapCNN does, and we did not yet have access to a Loihi2 edge device. The experimental setup is depicted in Fig. 4 on the right, and the physical setup is shown in the background of Fig. 1. In this experiment, we shot 15 table tennis balls with a Butterfly Amicus ball gun. Since the 3D triangulation and robot arm control are not part of this work, we used the existing ones from [8]. We report the ball return rate of the robot, since controlling the landing point of the ball is out of scope for this work. Given this setup, we achieved a ball return rate of $1.0$.
@@ V CONCLUSION
With neuromorphic hardware more accessible in recent years, Neuromorphic Computing (NC) and Spiking Neural Networks (SNNs) in particular has become more relevant for robotics. In this work, we used an event camera in combination with SNNs for ball detection as a real-time perception task. Three different SNN frameworks, namely sinabs, MetaTF, and Lava, and a corresponding neuromorphic edge device (DynapCNN, Akida, and Loihi2) were used, and their errors, time per forward pass, inference time, and power consumption were compared. Moreover, we show that an SNN on a neuromorphic hardware is able to run in a challenging table tennis robot setup in real-time.
Despite the promise of asynchronous processing with SNNs, the current edge devices face limitations, primarily attributed to hardware integration. Our results show that the better a neuromorphic edge device is connected to the main compute unit, e.g., as a PCIe card, the better the overall run-time.